Write the expression in standard form.

2/2-15i
1
Expert's answer
2014-06-05T07:52:54-0400
We need multiply and divide 2/(2-15i) by 2+15i. We obtain: 2(2+15i)/((2-15i)(2+15i))=(4+30i)/(4+225)=(4+30i)/229.

Answer. 2/(2-15i)=4/229+30i/229.